OpenStudy (anonymous):

Solve for a D=ut+0.5at^2

OpenStudy (australopithecus):

how far have you gotten through this problem?

OpenStudy (anonymous):

All the way to D-ut=0.5at^2

OpenStudy (australopithecus):

you know you can multiply or divide both sides of an equation by the same thing and not ruin the equality, just as you can add or subtract which is what you just did. Recall x/x = 1 3/3 = 1 1000/1000 = 1

OpenStudy (anonymous):

So im confused what do i do

OpenStudy (australopithecus):

you can multiply both sides of the equation by the same thing and it doesnt change the equality for example, a/5 = 3/c (a/5)*c = (3/c)*c <- key step ac/5 = 3c/c ac/5 = 3(1) ac/5 = 3 you can do the same with division

OpenStudy (australopithecus):

notice how I eliminated c from one side of the equation by simply multiplying both sides of the equality by c

OpenStudy (australopithecus):

do you know how you can solve your problem now?

OpenStudy (australopithecus):

Divide both sides of your equation by what you want to eliminate from the right side of the equation